OpenStudy (anonymous):

help me please! (1/6x2 ) = ( 1/3x2) - (1/x)

OpenStudy (anonymous):

\[\frac{ 1 }{ 6*2 }=\frac{1 }{ 3*2 }-\frac{ 1 }{ x }\]

OpenStudy (anonymous):

u need find a common denominator...

OpenStudy (anonymous):

\[\frac{ 1 }{ 12 }=\frac{ 1 }{ 6 }-\frac{ 1 }{ x}\]

OpenStudy (anonymous):

common denominator 12x so.... \[\frac{ 1(x) }{ 12(x) }=\frac{ 1(2x) }{ 6(2x) }-\frac{ 1(12) }{ x(12)}\]

OpenStudy (anonymous):

whatever you do to the bottom, u also have to do to the top^^^

OpenStudy (anonymous):

so u get... \[\frac{ 1x }{ 12x }=\frac{ 2x }{ 12x }-\frac{ 12 }{ 12x}\] u can get rid of the 12x on the bottom since it wont make a difference since its under all of the terms ... now we get 1x = 2x-12

OpenStudy (anonymous):

subtract 2x on both sides 1x =2x-12 -2x -2x u get: -1x=-12 Divide 1x on both sides, since its the term with the variable, and u want to get the variable alone.... so -1x=-12 /-1 /-1 final answer x=12

OpenStudy (anonymous):

FINAL ANSWER x=12,,, srry for taking up alot of space, lol